# Boosting Temporal Sentence Grounding via Causal Inference

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Information retrieval
@@ 2.2. Causal Inference
Causal inference (Pearl et al., 2016) including causal intervention and counterfactual reasoning has gained widespread application across various areas in the field of computer vision, including video question answer (Wang et al., 2020a; Niu et al., 2021), semantic segmentation (Zhang et al., 2020b; Chen et al., 2024), and image captioning (Liu et al., 2022b; Yang et al., 2023). It introduces a novel approach to deep learning, offering a way to identify the causal correlations between variables and mitigate spurious correlations. In TSG, DCM (Yang et al., 2021) leverages causal intervention to effectively disentangle and reduce location bias in videos. Simultaneously, IVG-DCL (Nan et al., 2021) applies causal intervention to mitigate biases inherent in language queries. However, both methods focus on a single modality, either video or textual query, and neglect the bias introduced by the other. This one-sided treatment limits their ability to fully address cross-modal spurious correlations. To tackle these issues, we propose CICR, which jointly tackles both textual and video biases through TCI and VCR. CAEM (Jiang et al., 2024a) simulates distribution shift via counterfactual sampling. However, it depends heavily on heuristic data augmentation, while our approach is more straightforward.
@@ 3.3. Textual Causal Intervention
As shown in Fig. 3(b), an unobserved confounder $Z$ introduces spurious correlations in the $Q\rightarrow M\rightarrow Y$ branch by affecting the likelihood $P(Y|V,Q)$. Traditional TSG models often directly learn the matching score $P(Y|V,Q)$ based on correlations between videos and queries. The non-interventional prediction can be expressed using the Bayes rule as:
@@
(2) $\begin{split}P(Y|V,do(Q))&=\sum_{z}P(Y|V,do(Q),z)P(z|V,do(Q))\\ &=\sum_{z}P(Y|V,Q,z)P(z|V,do(Q))\\ &=\sum_{z}P(Y|V,Q,z)P(z).\\ \end{split}$
@@
3.3.1. Confounder Dictionary Construction
A typical query in TSG usually consists of a subject, action, and object, aligning with the intuitive structure of “somebody doing something” in query descriptions(Nan et al., 2021; Liu et al., 2023a). To effectively capture these semantic relations, we adopt the Open Information Extraction (OpenIE) model (Stanovsky et al., 2018), which is well suited for extracting verb-centered relation tuples (i.e., subject, action, object) from large-scale natural language data. Based on extracted tuples, we construct three sets of vocabularies: a subject vocabulary $Z_{s}=\{z_{s}^{1},\ldots,z_{s}^{L_{s}}\}$, an action vocabulary $Z_{a}=\{z_{a}^{1},\ldots,z_{a}^{L_{a}}\}$, and an object vocabulary $Z_{o}=\{z_{o}^{1},\ldots,z_{o}^{L_{o}}\}$, where $L_{s}$, $L_{a}$, and $L_{o}$ indicate the respective vocabulary sizes. These components are then integrated into a confounder dictionary consisting of verb-centered relation vocabularies: $Z=[Z_{s};Z_{a};Z_{o}]=\{z_{1},z_{2},\ldots,z_{K}\}$, where $K=L_{s}+L_{a}+L_{o}$ represents the number of the total types of confounders. This structured representation serves as the foundation for modeling and mitigating textual biases introduced by annotator experience.
3.3.2. Backdoor Adjustment
Following (Wang et al., 2023a), we can adopt the Normalized Weighted Geometric Mean (NWGM) (Xu et al., 2015) to approximate the deconfounded prediction. The Eq. 2 can be formulated as follow:
@@ 3.4. Visual Counterfactual Reasoning
As such, we utilize the reliable indirect effect (i.e., TIE) for the unbiased prediction by counterfactual inference, as shown in Fig. 3(c). Note that $m^{*}$ denotes the situation where $v$ is not provided, i.e. $V=\varnothing$, which is void. As neural models cannot deal with empty inputs, we assume that the model will randomly guess with equal probability by setting $Y_{m^{*}}=c$ where $c$ denotes a learnable parameter.
@@ 4.1. Datasets and Evaluation Metrics
4.1.1. Datasets
1) TACoS includes much fewer but longer videos with many more sentences within a video, which is divided into three sets: training, validation, and testing, consisting of 10,146, 4589, and 4083 video-query pairs, respectively. 2) QVHighlights contains over 10,000 videos annotated with human-written text queries, serving as a fair benchmark since evaluation on the text split requires submission to the official server [^1] https://codalab.lisn.upsaclay.fr/competitions/6937. 3) Charades-CD is divided into four splits: training, validation, test-iid, and test-ood. The first three are independent and identically distributed, while the test-ood set out-of-distribution samples for evaluating debiasing. 4) Charades-CG consists of three testing splits: Test-Trivial, Novel-composition, and Novel-Word. Test-Trivial includes compositions seen during training, Novel-Composition features unseen combinations of known concepts, and Novel-Word contains words that do not appear in the training vocabulary.
4.1.2. Evaluation Metrics
Following previous work (Li et al., 2023b; Liu et al., 2024), we adopt R1@$m$, mIoU, mAP@$n$, and mAPavg as evaluation metrics. R1@$m$ measures the percentage of instances where the top-1 predicted moment exhibits an Intersection over Union (IoU) with the ground truth (GT) higher than $m$. mIoU reflects the average IoU with GT across all test samples. mAP@$n$ represents the mean average precision with an IoU greater than $n$, while mAPAvg is the average mAP@$n$ over IoU thresholds from 0.5 to 0.95 at 0.05 intervals.
@@ 4.3. Comparisons with the State-of-the-art
4.3.1. TACoS
We compare the performance of CICR with the SOTA methods, as shown in Tab. 1. The results demonstrate that CICR outperforms the SOTA methods and maintains top positions across all evaluation metrics. Specifically, when compared to the baseline MESM, CICR achieves an improvement of 0.35 in R1@0.3, 0.37 in R1@0.5, and 1.09 in mIoU, respectively. These improvements validate the effectiveness of our method.
4.3.2. QVHighlights
As shown in Table 2, our proposed CICR achieves superior performance compared to the baseline model MESM, with approximately 0.5 improvement across the R1@$n$ metrics. Although mAP@0.75 of CICR is slightly lower than the MESM by 0.06, it is worth noting that CICR surpasses MESM on more representative metric mAPAvg, achieving a gain of 0.77. Furthermore, CICR also yields a 1.21 improvement. These results demonstrate the effectiveness of CICR.
@@ 4.4. Experiments on Annotation Bias
4.4.1. Charades-CD
As shown in Tab. 3, our proposed CICR achieves the highest performance on most metrics on the Charades-CD. Specifically, CICR achieves an R1@0.3 of 70.73, R1@0.5 of 54.72 and the best mIoU of 47.89, outperforming the previous best baseline MESM by 1.04, 0.24, and 0.34 points, respectively. Although the stricter R1@0.7 score is slightly lower than MESM, the difference 0.09 is negligible. These results demonstrate that CICR exhibits strong generalization ability when encountering unseen or complex query-video patterns in OOD scenarios.
4.4.2. Charades-CG
Furthermore, to systematically investigate the model’s generalization capacity for infrequent linguistic constructs, we conducted experiments on Charades-CG. The results are shown in Tab. 4. Under Novel-Composition split, CICR achieves the best R1@0.5 and mIoU, outperforming both original and reproduced baseline MESM. While the R1@0.7 score is slightly lower than that of MESM, the mIoU improves that the total performance of CICR is better. For the Novel-Word split, CICR shows advantages by achieving the best performance across all metrics, indicating its stronger capability in generalizing to novel linguistic expressions.
@@ 4.5. Ablation Study
4.5.1. Effectiveness of Each Component
We conduct the ablation study of each element in our CICR framework on the QVHighlights val split. The key components of our CICR are the two causal modules: Textual Causal Intervention (TCI) and Visual Counterfactual Reasoning (VCR). As shown in Tab. 5, each component is beneficial for TSG. TCI and VCR achieve gains of 0.74% and 0.29% in mAPAvg., respectively. When we use all of them, the result comes to the best, which gains 1.36% in mAP@0.75 and 1.36% in mAPAvg., demonstrating that both of them are beneficial to the results.
4.5.2. Effectiveness of loss $\mathcal{L}_{kl}$
As shown in Fig. 5, we conduct an ablation study to investigate the contribution of the KL divergence loss $\mathcal{L}_{kl}$ in CICR. When $\mathcal{L}_{kl}$ is removed, we observe a performance drop across all evaluation metrics. This highlights the crucial role of the KL loss in ensuring the reliable estimation of counterfactual situation $Y_{m^{*}}$.
4.5.3. Hyper-parameter Sensitive Analysis
The hyper-parameter $\lambda_{kl}$ controls the balance of the KL loss $\mathcal{L}_{kl}$ and TSG loss $\mathcal{L}_{tsg}$. In Fig. 6, we investigate the impact of varying values of $\lambda_{kl}$ on model performance. It is shown that for the TACoS, the best mIoU is achieved when $\lambda_{kl}=0.05$. For the QVHighlights, the highest mAPAvg is obtained with $\lambda_{kl}=0.1$.
4.5.4. Generality of the method
To further validate the generality and flexibility of our proposed method, we integrate CICR into QD-DETR (Moon et al., 2023) and evaluate its performance on the QVHighlights val split. As shown in Tab. 6, the incorporation of CICR consistently leads to performance improvements across all metrics, demonstrating that our debiasing strategy is not limited to a specific architecture. This result confirms that CICR can serve as a plug-and-play module that effectively enhances various existing temporal grounding models by mitigating modality-induced biases and improving generalization.
@@ 5. Conclusion
6. Acknowledgments
This work is supported in part by the National Natural Science Foundation of China under Grant xxxx.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755624isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Activity recognition and understanding
@@ 2.2. Causal Inference
Causal inference (Pearl et al., 2016) including causal intervention and counterfactual reasoning has gained widespread application across various areas in the field of computer vision, including video question answer (Wang et al., 2020a; Niu et al., 2021), semantic segmentation (Zhang et al., 2020b; Chen et al., 2024a), and image captioning (Liu et al., 2022b; Yang et al., 2023). It introduces a novel approach to deep learning, offering a way to identify the causal correlations between variables and mitigate spurious correlations. In TSG, DCM (Yang et al., 2021) leverages causal intervention to effectively disentangle and reduce location bias in videos. Simultaneously, IVG-DCL (Nan et al., 2021) applies causal intervention to mitigate biases inherent in language queries. However, both methods focus on a single modality, either video or textual query, and neglect the bias introduced by the other. This one-sided treatment limits their ability to fully address cross-modal spurious correlations. To tackle these issues, we propose CICR, which jointly tackles both textual and video biases through TCI and VCR. CAEM (Jiang et al., 2024a) simulates distribution shift via counterfactual sampling. However, it depends heavily on heuristic data augmentation, while our approach is more straightforward.
@@ 3.3. Textual Causal Intervention
As shown in Fig. 3(b), an unobserved confounder $Z$ introduces spurious correlations in the $Q\rightarrow M\rightarrow Y$ branch by affecting the likelihood $P(Y|V,Q)$. This latent confounding may lead the model to rely on superficial statistical associations rather than true semantic alignment. Traditional TSG models often directly learn the matching score $P(Y|V,Q)$ based on correlations between videos and queries. The non-interventional prediction can be expressed using the Bayes rule as:
@@
(2) $\begin{split}P(Y|V,do(Q))&=\sum_{z}P(Y|V,do(Q),z)P(z|V,do(Q))\\ &=\sum_{z}P(Y|V,Q,z)P(z).\end{split}$
@@
Confounder Dictionary Construction. A typical query in TSG usually consists of a subject, action, and object, aligning with the intuitive structure of “somebody doing something” in query descriptions(Nan et al., 2021; Liu et al., 2023a). To effectively capture these semantic relations, we adopt the Open Information Extraction (OpenIE) model (Stanovsky et al., 2018), which is well suited for extracting verb-centered relation tuples (i.e., subject, action, object) from large-scale natural language data. Based on extracted tuples, we construct three sets of vocabularies: a subject vocabulary $Z_{s}=\{z_{s}^{1},\ldots,z_{s}^{L_{s}}\}$, an action vocabulary $Z_{a}=\{z_{a}^{1},\ldots,z_{a}^{L_{a}}\}$, and an object vocabulary $Z_{o}=\{z_{o}^{1},\ldots,z_{o}^{L_{o}}\}$, where $L_{s}$, $L_{a}$, and $L_{o}$ indicate the respective vocabulary sizes. These components are then integrated into a confounder dictionary consisting of verb-centered relation vocabularies: $Z=[Z_{s};Z_{a};Z_{o}]=\{z_{1},z_{2},\ldots,z_{K}\}$, where $K=L_{s}+L_{a}+L_{o}$ represents the number of the total types of confounders. This structured representation serves as the foundation for modeling and mitigating textual biases introduced by annotator experience.
Backdoor Adjustment. Following (Wang et al., 2023a; Chen et al., 2024b), we can adopt the Normalized Weighted Geometric Mean (NWGM) (Xu et al., 2015) to approximate the deconfounded prediction. The Eq. 2 can be formulated as follow:
@@ 3.4. Visual Counterfactual Reasoning
Thus, we utilize the reliable indirect effect (i.e., TIE) for the unbiased prediction by counterfactual inference, as shown in Fig. 3(c). Note that $m^{*}$ denotes the situation where $v$ is not provided, i.e. $V=\varnothing$, which is void. As neural models cannot deal with empty inputs, we assume that the model will randomly guess with equal probability by setting $Y_{m^{*}}=c$ where $c$ denotes a learnable parameter.
@@ 4.1. Datasets and Evaluation Metrics
Datasets. 1) TACoS includes much fewer but longer videos with many more sentences within a video, which is divided into three sets: training, validation, and testing, consisting of 10,146, 4589, and 4083 video-query pairs, respectively. 2) QVHighlights contains over 10,000 videos annotated with human-written text queries, serving as a fair benchmark since evaluation on the text split requires submission to the official server [^1] https://codalab.lisn.upsaclay.fr/competitions/6937. 3) Charades-CD is divided into four splits: training, validation, test-iid, and test-ood. The first three are independent and identically distributed, while the test-ood set out-of-distribution samples for evaluating debiasing. 4) Charades-CG consists of three testing splits: Test-Trivial, Novel-composition, and Novel-Word. Test-Trivial includes compositions seen during training, Novel-Composition features unseen combinations of known concepts, and Novel-Word contains words that do not appear in the training vocabulary.
Evaluation Metrics. Following previous work (Li et al., 2023b; Liu et al., 2024), we adopt R1@$m$, mIoU, mAP@$n$, and mAPavg as evaluation metrics. R1@$m$ measures the percentage of instances where the top-1 predicted moment exhibits an Intersection over Union (IoU) with the ground truth (GT) higher than $m$. mIoU reflects the average IoU with GT across all test samples. mAP@$n$ represents the mean average precision with an IoU greater than $n$, while mAPAvg is the average mAP@$n$ over IoU thresholds from 0.5 to 0.95 at 0.05 intervals.
@@ 4.3. Comparisons with the State-of-the-art
TACoS. We compare the performance of CICR with the SOTA methods, as shown in Tab. 1. The results demonstrate that CICR outperforms the SOTA methods and maintains top positions across all evaluation metrics. Specifically, when compared to the baseline MESM, CICR achieves an improvement of 0.35 in R1@0.3, 0.37 in R1@0.5, and 1.09 in mIoU, respectively. These improvements validate the effectiveness of our method.
QVHighlights. As shown in Table 2, our proposed CICR achieves superior performance compared to the baseline model MESM, with approximately 0.5 improvement across the R1@$n$ metrics. Although mAP@0.75 of CICR is slightly lower than the MESM by 0.06, it is worth noting that CICR surpasses MESM on more representative metric mAPAvg, achieving a gain of 0.77. Furthermore, CICR also yields a 1.21 improvement. These results demonstrate the effectiveness of CICR.
@@ 4.4. Experiments on Annotation Bias
Charades-CD. As shown in Tab. 3, our proposed CICR achieves the highest performance on most metrics on the Charades-CD. Specifically, CICR achieves an R1@0.3 of 70.73, R1@0.5 of 54.72 and the best mIoU of 47.89, outperforming the previous best baseline MESM by 1.04, 0.24, and 0.34 points, respectively. Although the stricter R1@0.7 score is slightly lower than MESM, the difference 0.09 is negligible. These results demonstrate that CICR exhibits strong generalization ability when encountering unseen or complex query-video patterns in OOD scenarios.
Charades-CG. Furthermore, to systematically investigate the model’s generalization capacity for infrequent linguistic constructs, we conducted experiments on Charades-CG. The results are shown in Tab. 4. Under Novel-Composition split, CICR achieves the best R1@0.5 and mIoU, outperforming both original and reproduced baseline MESM. While the R1@0.7 score is slightly lower than that of MESM, the mIoU improves that the total performance of CICR is better. For the Novel-Word split, CICR shows advantages by achieving the best performance across all metrics, indicating its stronger capability in generalizing to novel linguistic expressions.
@@
Table 7. Efficiency comparisons with the baseline MESM on the QVHighlights dataset.
Methods Params (M) Epoch Time (sec)
MESM 14.13 73
CICR (ours) 14.27 75
@@ 4.5. Ablation Study
Effectiveness of Each Component. We conduct the ablation study of each element in our CICR framework on the QVHighlights val split. The key components of our CICR are the two causal modules: Textual Causal Intervention (TCI) and Visual Counterfactual Reasoning (VCR). As shown in Tab. 5, each component is beneficial for TSG. TCI and VCR achieve gains of 0.74% and 0.29% in mAPAvg., respectively. When we use all of them, the result comes to the best, which gains 1.36% in mAP@0.75 and 1.36% in mAPAvg., demonstrating that both of them are beneficial to the results.
Effectiveness of loss $\mathcal{L}_{kl}$. As shown in Fig. 5, we conduct an ablation study to investigate the contribution of the KL divergence loss $\mathcal{L}_{kl}$ in CICR. When $\mathcal{L}_{kl}$ is removed, we observe a performance drop across all evaluation metrics. This highlights the crucial role of the KL loss in ensuring the reliable estimation of counterfactual situation $Y_{m^{*}}$.
Hyper-parameter Sensitive Analysis. The hyper-parameter $\lambda_{kl}$ controls the balance of the KL loss $\mathcal{L}_{kl}$ and TSG loss $\mathcal{L}_{tsg}$. In Fig. 6, we investigate the impact of varying values of $\lambda_{kl}$ on model performance. It is shown that for the TACoS, the best mIoU is achieved when $\lambda_{kl}=0.05$. For the QVHighlights, the highest mAPAvg is obtained with $\lambda_{kl}=0.1$.
Generality of the method. To validate the generality and flexibility of CICR, we integrate it into QD-DETR (Moon et al., 2023) and evaluate on the QVHighlights val split. As shown in Tab. 6, CICR consistently improves performance across all metrics, demonstrating that our debiasing strategy is not limited to a specific architecture. This result confirms that CICR can serve as a plug-and-play module that effectively enhances various existing temporal grounding models by mitigating modality-induced biases and improving generalization.
Computational complexities. We assess the computational complexity of CICR against the baseline MESM by parameter count and epoch time on the QVHighlights dataset using an NVIDIA RTX 3090 Ti GPU. As shown in Tab. 7, CICR adds negligible parameters and maintains comparable epoch time, demonstrating improved performance with minimal overhead.
@@ 5. Conclusion
Acknowledgements.
This research was supported partially by the National Natural Science Foundation of China (Grant Nos. 62276203, 62036007) and the Double First-Class Overseas Research Project of Xidian University.